OpenStudy (anonymous):

A tugboat goes upstream 160 miles in 10 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current.

OpenStudy (lgbasallote):

upstream speed = speed of boat - speed of current downstream speed = speed of boat + speed of current it says upstream speed is 160 miles in 10 hours so that's \(\frac{160 \; \text{miles}}{10 \; \text{hours}} = 16 \; \text{mph}\) downstream speed is 160 miles in 5 hours so 160/5 = 32 mph now let speed of boat = b speed of current = c upstream speed = b - c downstream speed = b+ c 16 = b- c 32 = b+c you're looking for speed wihtout current so solve for b (speed of boat)